OpenStudy (anonymous):

A tin can (right circular cylinder) with top and bottom is to have volume V . What dimensions (the radius of the bottom and the height) give the minimum total surface area?

OpenStudy (anonymous):

V=h(pi)r^2 , A=h2(pi)r + 2(pi)r^2

OpenStudy (anonymous):

solving for h, h=V/((pi)r^2)

OpenStudy (anonymous):

\[A=2V/r + 2 \pi r^2\]

OpenStudy (anonymous):

Does that look right?

OpenStudy (phi):

so far, so good. I assume this is a calculus problem ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now, I take the derivative of that, and set it equal to zero?

OpenStudy (phi):

yes

OpenStudy (anonymous):

what about still having two variables on RHS?

OpenStudy (phi):

in this problem, volume is "given" and a fixed value I would use the notation \(V_0\) to emphasize it is a constant

OpenStudy (anonymous):

so, the derivative of V is 0?

OpenStudy (phi):

if we had to find its derivative, it would be 0 but in your equation, it is a coefficient.

OpenStudy (phi):

you want to find dA/dr of \[ \frac{dA}{dr}= \frac{d}{dr}\left(2V_0\ r^{-1} + 2 \pi r^2 \right)\]

OpenStudy (anonymous):

\[0 = -2Vr^-2 + 4 \pi r\]

OpenStudy (phi):

yes

OpenStudy (phi):

any luck with this ?

OpenStudy (anonymous):

sorry, had to step away

OpenStudy (anonymous):

came up with \[r=\sqrt[3]{V/2 \pi}\]

OpenStudy (anonymous):

which equation do I sub r into to find h